# Training-free Task-oriented Grasp Generation

### v1

@@ Abstract
This paper presents a training-free pipeline for task-oriented grasp generation that combines pre-trained grasp generation models with vision-language models (VLMs). Unlike traditional approaches that focus solely on stable grasps, our method incorporates task-specific requirements by leveraging the semantic reasoning capabilities of VLMs. We evaluate five querying strategies, each utilizing different visual representations of candidate grasps, and demonstrate significant improvements over a baseline method in both grasp success and task compliance rates, with absolute gains of up to 36.9% in overall success rate. Our results underline the potential of VLMs to enhance task-oriented manipulation, providing insights for future research in robotic grasping and human-robot interaction.
@@ I Introduction
Several grasp generation methods have been introduced in recent research [8, 17, 15, 21], with most focusing on achieving stable, collision-free grasps. While stability and collision avoidance are critical, human grasping behavior is inherently task-driven. For humans, grasping is not just about securing an object safely but about fulfilling a specific goal. For instance, when picking up a cup to drink, a person typically grasps it by the handle, even though other stable, collision-free grasps — like holding the cup’s body — are possible. This task-oriented approach highlights that grasp selection is based on the intended use of the object—an idea not fully embraced by traditional robotic grasping techniques.
@@
In contrast, vision-language models (VLMs) [29, 12], which are trained on massive, internet-sourced datasets, have demonstrated remarkable generalization capabilities in understanding both natural language and images. This raises an intriguing question: can VLMs be leveraged to aid in task-oriented grasp generation? Several studies have explored this potential [1, 23]. However, these approaches often require additional training steps to generate task-oriented grasps (see Section II) and tend to be tightly integrated with a particular grasp generation model. This rigidity limits flexibility and adaptability, making it challenging to incorporate more advanced grasp generation models as they emerge.
In this project, we propose a training-free pipeline for task-oriented grasp generation by combining a pre-trained grasp generation model with a vision-language model (VLM). The key insight is that the pre-trained grasp generation model can be augmented to produce task-oriented grasps by leveraging a VLM to filter or select candidate grasps based on task-specific requirements.
We extensively evaluated and compared different approaches for integrating grasp generation models with VLMs in a simulation environment [9]. Our experiments demonstrate that this simple yet effective strategy achieves a 36.9% absolute improvement in the overall success rate. This study provides valuable insights into how pre-trained models and VLMs can be seamlessly integrated to enhance task-oriented grasping.
@@
This report is structured as follows:
  • Section II reviews related work on grasp generation and foundation models in robotics.
  • Section III defines the problem and presents our proposed solution.
  • Section IV describes the experimental setup, evaluation methods, and results.
  • Section V concludes the report and outlines directions for future research.
Figure 1: Overview of the system. The depth image is unprojected into a point cloud, which is processed by a grasp generation model (e.g., Contact GraspNet) to produce a set of unconditional grasps. Simultaneously, an open-vocabulary object detector identifies the target object from the RGB image, generating a segmentation mask that filters the grasps. The top K grasps are selected based on confidence and further refined by a motion planner to ensure trajectory feasibility. These filtered grasps are then evaluated using a vision-language model (VLM) to select the best grasp for the task.
Figure 2: Illustration of the five query methods used for grasp evaluation. The images sent to the VLM vary based on the query method, differing in how the grasps are visualized: represented either by a fork-shaped mesh or a colored dot, displayed as individual images (one grasp per image) or combined into a single image (all grasps in one image). For the CPG method, only the raw image is sent, and the VLM is asked to indicate the grasping point.
@@ III-B System Overview
Our proposed approach decomposes the estimation of $P(G^{*}\mid X,T)$ into two stages: (1) task-agnostic grasp sampling $P(G^{*}\mid X)$, and (2) task-oriented grasp evaluation $P(S\mid X,T,g)$. This factorization allows us to leverage existing work focused on stable grasp generation while incorporating task-oriented constraints to ensure the grasp is suitable for the task at hand. An overview of our proposed system is illustrated in Fig. 1.
@@
Simultaneously, we employ an open-vocabulary object detector [16] to identify the target object from the RGB image, generating a segmentation mask for the detected object. This mask is used to filter the candidate grasps, retaining only those associated with the target object.
@@
Generated grasps are passed to a motion planner, which filters out any grasps that are not feasible due to reachability or collision constraints. This step is performed prior to querying the vision-language model (VLM) to reduce computational overhead.
Figure 3: Illustration of using K-means clustering and scores to select diverse grasps while taking into account for the quality of each grasp.
III-C Grasp Clustering
@@
One limitation of the proposed pipeline is that the pre-filtered top-$k$ grasp candidates (based on the confidence scores from the grasp generation model) often lack diversity. These candidates typically belong to the same cluster of grasps. For instance, in the case of a screwdriver, most top-rated grasps focus on the midpoint of the handle (Fig. 4). While these grasps yield a high grasp success rate, they are less useful for task-oriented grasp generation, which requires a diverse set of grasps that can then be ranked by the Vision-Language Model (VLM).
@@
To address this issue, we generate $N$ valid grasps (pre-filtered by the motion planner as described earlier in section III-B) and perform $k$-means clustering on the contact points (i.e., the center of the gripper) in Euclidean space. For our experiments, we set $K=3$. From each cluster, we select the grasp with the highest confidence score as predicted by the grasp generation model. This strategy ensures that the final set of candidate grasps $\{\bar{G}_{1},\bar{G}_{2},\bar{G}_{3}\}$ is both diverse and of high quality (Fig. 3).
@@
It is worth noting that more sophisticated clustering methods could also be employed. For example, $k$-means clustering could be applied to features extracted from a vision foundation model such as DINO [30]. In this case, each cluster would represent grasps that target a similar part of the object, while different clusters would correspond to grasps targeting distinct parts of the object. However, in this work, we use the simpler strategy outlined above, as it empirically yields satisfactory results.
III-D Task-Oriented Grasp Filtering by VLM
@@
After generating $K=3$ grasp candidates, we query the Vision-Language Model (VLM) to select the grasp most suitable for the specified task. Queries are constructed by combining the grasp candidates with the task description and submitting them to the VLM. The VLM’s response is then used to select the best grasp for the task.
@@
We employed five distinct strategies to query the VLM, each differing in how the candidate grasps are visualized and presented:
@@
    Grasps in a Single Image (GSI): All candidate grasps are visualized in a single image, with each grasp represented by the gripper shape at its corresponding pose on the object’s point cloud. This composite visualization enables the VLM to compare all grasps directly within one scene.
@@
    Contact Points in a Single Image (CPSI): All candidate grasps are visualized as their respective contact points (colored dots) in a single image. The VLM selects the best grasp based on the task description, facilitating direct comparison of contact points in one image.
@@
    Grasps in Multiple Images (GMI): Each candidate grasp is presented in a separate image, with the grasp pose depicted by a simplified gripper shape positioned on the object’s point cloud. The gripper mesh, generated using the trimesh [6] library, is designed for clarity. The VLM selects the image that contains the grasp best suited for the task.
@@
    Contact Points in Multiple Images (CPMI): Each candidate grasp is visualized in a separate image, represented by a colored dot indicating the center of the grasp on an RGB image. This simplified visualization aims to reduce visual complexity for the VLM.
@@
    Contact Point Generation (CPG): Instead of providing predefined grasp candidates, we directly query the VLM to suggest a grasping point on the image that best satisfies the task description by specifying pixel coordinates $(u,v)$. To ensure the selected point can be unprojected into 3D space, we explicitly request that the VLM identify a point on the target object. We then choose the 3D point $P$ whose projection onto the image plane is closest to the $(u,v)$ coordinate. Subsequently, we select the grasp whose contact point is nearest to the 3D point $P$.
An illustration of these query methods is provided in Fig. 2.
@@
TABLE I: Experiment results of the proposed pipeline using different VLM query strategies for each VLM.
IV Experiments
We conducted experiments to evaluate our training-free, task-oriented grasp generation pipeline using various vision-language models (VLMs) and querying strategies. Our goal was to assess the effectiveness of integrating pre-trained grasp generation models with VLMs to select grasps that satisfy specific task requirements.
IV-A Experimental Setup
Our experiments were performed in the ManiSkill simulator [9], using the Franka Emika Panda robotic arm equipped with a parallel-jaw gripper. We selected 29 objects from the YCB [4] and Google Scanned Objects (GSO) [2] datasets, ensuring a diverse range of shapes, sizes, and functionalities that align with our predefined tasks. These objects were chosen to fit the specific task scenarios we designed.
For each object, we defined a set of tasks with specific grasping requirements. We formulated five broad task categories, such as “gripping a specific part of a given tool” and “gripping a specific part of a given toy animal.” Each category was further divided into 2 or 3 sub-tasks, resulting in a total of 14 unique task descriptions. Examples include “gripping the tail of the toy animal” and “gripping the head of the toy animal” which form a pair of sub-tasks, and “gripping the handle of the tool to ensure a secure hold” and “gripping the tool by a part other than the handle for handing over”, forming another pair of sub-tasks under one of the 5 tasks. (For more details on our definition of the tasks, please refer to Appendix A VI.)
IV-B Grasp Generation and Selection
We utilized Contact GraspNet (CGN) [21] to generate a set of candidate grasps for each object. CGN provides grasps along with associated quality scores. A motion planner (mplib) [14] was applied to filter out infeasible grasps due to reachability issues or potential collisions. We then followed the steps described in Sections III-B and III-C to generate a final set of diverse, high-quality grasp candidates.
We employed five strategies for querying the VLM, as detailed in Section III-D, which differ in how the candidate grasps are visualized and presented. Five corresponding prompts were carefully crafted (fig 2) to ensure the VLM could effectively interpret the task, reason through the grasp selection process, and provide responses in a consistent and easily parsable format. (For details on the prompts, please refer to Appendix B.)
We experimented with two VLMs:
  • Qwen-7B [26]: A state-of-the-art vision-language model known for its strong language understanding and visual reasoning capabilities.
  • Molmo [7]: A vision-language model designed specifically for robotic applications, including point-based interactions, making it particularly suitable for the PSI strategy.
For each VLM, we applied all five query strategies, resulting in a total of ten configurations. While a more powerful VLM (e.g., Qwen-72B) could potentially enhance performance, we used the 7B versions for both models due to computational constraints.
We compared the performance of these ten configurations against a baseline system, which selects the grasp generated by the grasp generation model with the highest confidence score. This baseline does not involve task-specific reasoning or diversity in grasp selection, serving as a straightforward benchmark for evaluating the advantages of the proposed pipeline.
IV-C Evaluation Metrics
@@
IV-D Automated Evaluation
Manual evaluation of task compliance by humans can be time-consuming. To automate this process, we used a more powerful VLM, Qwen-72B [26], as an evaluator. This model received the task description, the selected grasp, and the three candidate grasps, then provided an explanation followed by a final judgment of ”Yes” or ”No” indicating task compliance. The explanations allowed for improved accuracy and traceability of the judgments. (Details on the prompts used for querying Qwen-72B for automated evaluation are provided in Appendix C.)
To validate the reliability of this automated evaluation, we manually reviewed a subset of the results and compared them against Qwen-72B’s judgments. The precision of the automated evaluation (i.e., the agreement between Qwen-72B’s predictions and human evaluations) is 0.717.
The main limitation observed is that while humans rely on videos to evaluate the results, the Qwen-72B model bases its predictions solely on static images highlighting the grasp point. For certain objects, such as water bottles, a top-down view of the grasp point often fails to convey whether the grasp is targeting the handle or another part of the object, leading to ambiguity. To address this, we propose either utilizing a more powerful model or providing more informative images, such as capturing the object from multiple viewpoints.
We observe that the Qwen-72B model achieves a high task compliance prediction accuracy ($71.69\%$), though it is not entirely free from errors. Most prediction inaccuracies stem from the aforementioned ambiguities. Consequently, we chose to rely on human evaluations for the final results. We hypothesize that adopting more advanced models (e.g., GPT-4o) or incorporating richer visual information could significantly enhance task compliance prediction, potentially achieving reliability suitable for fully automated evaluation.
IV-E Experiment Results
Table I summarizes our experimental results. Notably, our proposed system consistently outperforms the baseline system, achieving an absolute improvement in combined success rate ranging from 17.5% to 54.4%, depending on the query strategy. Among these strategies, the Contact Point Generation (CPG) method demonstrates the best performance, surpassing the baseline by 36.9% in combined success rate. While the CPG method excels in reasoning tasks, gripper shape methods (GSI, GMI) show better grasp success rates, and contact point methods (CPSI, CPMI) perform better in task compliance, reflecting their alignment with task-specific requirements. This demonstrates the importance of tailoring visual and textual strategies to different model strengths for optimal performance.
We observe that the grasp success rate for methods utilizing the VLM is comparable to that of the baseline method, indicating that incorporating the VLM does not compromise grasp quality. Furthermore, the task compliance rate for methods leveraging the VLM is significantly higher than the baseline, demonstrating the effectiveness of our proposed system in aligning grasps with task-specific requirements.
Single image vs multiple images The results highlight the distinction between single-image and multi-image strategies in grasp evaluation. Single-image approaches, such as GSI and CPSI, generally perform better in task compliance, possibly due to the simplicity of presenting all grasps in one visual context. In contrast, multi-image strategies, like GMI and CPMI, spread grasps across multiple images, which may complicate the VLM’s ability to effectively compare them. For Qwen and Molmo, single-image methods achieved higher combined success rates, suggesting they better leverage the VLM’s ability to process integrated visual and textual information.
Gripper shape or contact point These results clearly reflect the difference in performance with respect to gripper shape versus contact point methods for the evaluation of grasps. Gripper shape methods (GSI and GMI) consistently achieve higher grasp success rates but lag in task compliance rates compared to contact point methods (CPSI and CPMI). This suggests that gripper shape representations provide better spatial understanding for stable grasps, while contact point methods, which simplify visual complexity, are more effective at aligning with task-specific requirements. Notably, CPSI achieves the highest task compliance rate with Molmo, while CPMI performs best in combined success rate, highlighting the importance of tailored visual representation for task-oriented grasps.
Comparison of VLM In the comparison of the two Vision-Language Models (VLMs), Qwen-7B demonstrates better performance in the Contact Point Generation (CPG) method, where no candidate grasps are explicitly provided. This suggests that Qwen excels at reasoning directly about task-oriented grasping points based solely on spatial and semantic cues without relying on visualized candidates. In contrast, Molmo achieves higher combined success rates in methods that provide candidate visualizations, such as CPSI and CPMI, indicating its strength in interpreting explicit visual grasp candidates. These observations emphasize that while Qwen-7B performs exceptionally well in reasoning tasks, Molmo benefits more from structured and simplified candidate representations for task-oriented grasp evaluation.
Figure 4: Example where clustering fails to produce diverse grasp candidates, resulting in grasps concentrated in similar regions of the object (at the handle of the screwdriver).
Figure 5: Example of an ambiguous case involving tasking the VLM to select a grasp that grips the lid/body of the bottle while the image given is in top-down view, making it hard for VLM to interpret different grasps.
IV-F Failure Modes

### v2

# TOG: Training-free Task-oriented Grasp Generation

@@ Abstract
We presents a training-free pipeline for task-oriented grasp generation (TOG) that integrates pre-trained grasp generation models with vision-language models (VLMs). Beyond focusing on stability, our method incorporates task-specific requirements by leveraging the semantic reasoning capabilities of VLMs. We evaluate different strategies for integrating the VLM with a pre-trained grasp generation model, demonstrating through extensive empirical studies that this simple yet effective approach performs robustly in both simulated environments and real-world robotic experiments. Our results highlight the potential of VLMs to enhance task-oriented manipulation, offering valuable insights for future research in robotic grasping and human-robot interaction. Codes can be found at: https://github.com/StevenLiudw/TOG
@@ I Introduction
Several grasp generation methods have been introduced in recent research [2, 7, 4, 3], with the majority focusing on achieving stable, collision-free grasps. While stability and collision avoidance are critical, human grasping behavior is inherently task-driven. For humans, grasping is not just about securing an object safely but about fulfilling a specific goal. For instance, when picking up a cup to drink, a person typically grasps it by the handle, even though other stable, collision-free grasps — like holding the cup’s body — are possible. This task-oriented approach highlights that grasp selection is based on the intended use of the object — an idea not fully embraced by traditional robotic grasping techniques.
@@
Figure 1: Given a specific task, TOG can generate feasible grasps according to the task specifications by leveraging the combined capabilities of Contact Grasp Net (CGN) and a Vision-Language Model (VLM).
In contrast, vision language models (VLMs) [8, 9], which are trained on massive, internet-scale datasets, have demonstrated remarkable generalization capabilities in understanding both natural language and images. This raises an intriguing question: can VLMs be leveraged to aid in task-oriented grasp generation? Several studies have explored this potential [11, 10]. However, these approaches often require additional training steps to generate task-oriented grasps (see Section II) and tend to be tightly integrated with a particular grasp generation model. This rigidity limits flexibility and adaptability, making it challenging to incorporate more advanced grasp generation models as they emerge.
In this project, we propose a training-free pipeline for task-oriented grasp generation (TOG) by integrating a pre-trained grasp generation model with a VLM. Our key insight is that for more capable VLMs, allowing them to freely specify grasp points and match them to generated grasps yields the best performance. In contrast, for less powerful models, a more constrained approach — where the VLM selects from pre-filtered grasps — leads to better results.
Figure 2: Overview of the system. The depth image is unprojected into a point cloud, which is processed by a grasp generation model (e.g., Contact GraspNet) to produce a set of unconditional grasps. The top K grasps are selected based on confidence and further refined by a motion planner to ensure trajectory feasibility. These filtered grasps are then evaluated using a vision-language model (VLM) to select the best grasp for the task.
We extensively evaluated and compared different approaches to integrating grasp generation models with VLMs in both a simulation environment [17] and real robot settings. Our experiments demonstrate that this simple yet effective strategy is highly robust, achieving an average 70% success rate on real robots. This study provides valuable insights into how pre-trained models and VLMs can be seamlessly integrated to enhance task-oriented grasping.
@@ II-C Foundation Models in Robotics
Figure 3: Illustration of using K-means clustering and scores to select diverse grasps while taking into account the quality of each grasp.
@@ III-B System Overview
Our proposed approach generates the optimal task-oriented grasp in two stages: (1) sampling a set of task-agnostic grasps from a pretrained grasp generation model, and (2) filtering and selecting the grasp that best satisfies the task constraints. This decomposition enables us to leverage existing work on stable grasp generation while integrating task-specific constraints to ensure the selected grasp is well-suited for the given task. An overview of our proposed system is illustrated in Fig. 2.
@@
Generated grasps are passed to a motion planner, which filters out any grasps that are not feasible due to reachability or collision constraints. This step is performed prior to querying the VLM to reduce computational overhead.
Image Prompt Text Prompt
Contact Points in a Single Image (CPSI)
[Uncaptioned image] Given an image showing the object with different grasp candidates highlighted in red, green, and blue dots, your task is to analyze these grasp candidates based on the task description and choose the index of the grasp that would best accomplish the task.
Contact Points in Multiple Images (CPMI)
[Uncaptioned image] Given multiple images, each showing the object with a different grasp candidate with a red dot, and a task description, your task is to analyze the grasp candidates in each image based on the task description and choose the index of the grasp that would best accomplish the task.
Grasps in a Single Image (GSI)
[Uncaptioned image] Given an image showing the object with different grasp candidates highlighted in red, green, and blue, your task is to analyze these grasp candidates based on the task description and choose the index of the grasp that would best accomplish the task.
Grasps in Multiple Images (GMI)
[Uncaptioned image] Given multiple images, each showing the object with a different grasp candidate highlighted in red, and a task description, your task is to analyze the grasp candidates in each image based on the task description and choose the index of the grasp that would best accomplish the task.
Contact Point Generation (CPG)
[Uncaptioned image] Given an image of the scene and a task description, you should provide a point on the image where the robot should grasp the object to best accomplish the task.
TABLE I: Illustration of different querying methods for VLM.
@@
One limitation of the proposed pipeline is that the pre-filtered top-$k$ grasp candidates (based on the confidence scores from the grasp generation model) often lack diversity. These candidates typically belong to the same cluster of grasps. For instance, in the case of a screwdriver, most top-rated grasps focus on the midpoint of the handle (Fig. 5 (a)). While these grasps yield a high grasp success rate, they are less useful for task-oriented grasp generation, which requires a diverse set of grasps that can then be ranked by the VLM.
@@
To enhance grasp diversity, we generate N valid grasps and pre-filter them using the motion planner, as described in Section III-B. We then apply $k$-means clustering to the contact points (i.e., the center of the gripper) in Euclidean space. The effect of different $k$ values is analyzed in Section IV-C. From each cluster, we select the grasp with the highest confidence score predicted by the grasp generation model. This approach ensures that the final set of candidate grasps is both diverse and of high quality (Fig. 3).
@@
It is worth noting that more sophisticated clustering methods could also be employed. However, in this work, we use the simpler strategy outlined above, as it empirically yields satisfactory results.
III-C Task-Oriented Grasp Filtering by VLM
@@
After generating K grasp candidates, we query the VLM to select the grasp most suitable for the specified task. Queries are constructed by combining the grasp candidates with the task description and submitting them to the VLM. The VLM’s response is then used to select the best grasp for the task.
@@
We employ five distinct strategies to query the VLM, each differing in how candidate grasps are visualized and presented.
@@
    Grasps in a Single Image (GSI): All candidate grasps are visualized in a single image, with each grasp represented as a U-shaped wireframe at its corresponding pose, projected onto the observed RGB image. This composite visualization allows the VLM to compare all grasps directly within a single scene.
@@
    Contact Points in a Single Image (CPSI): Candidate grasps are visualized as colored dots representing their contact points in a single image. The VLM selects the optimal grasp based on the task description, enabling direct comparison of contact points within a single view.
@@
    Grasps in Multiple Images (GMI): Each candidate grasp is displayed in a separate image, where the grasp pose is represented by a gripper wireframe projected onto an individual RGB observation. The VLM then selects the image containing the grasp best suited for the task.
@@
    Contact Points in Multiple Images (CPMI): Each candidate grasp is represented by a colored dot indicating its contact point, displayed in a separate image. This simplified visualization reduces visual complexity for the VLM, allowing it to focus solely on contact point selection.
@@
    Contact Point Generation (CPG): Instead of providing predefined grasp candidates, we directly query the VLM to suggest an optimal grasping point in the image by specifying pixel coordinates $(u,v)$. To ensure the selected point can be unprojected into 3D space, we explicitly instruct the VLM to identify a point on the target object. We then select the 3D point $P$ whose projection onto the image plane is closest to the queried $(u,v)$ coordinate. Finally, we choose the grasp whose contact point is nearest to the 3D point $P$.
An illustration of these query methods is provided in Table I.
IV Simulation Experiment
We conducted experiments in the simulator to evaluate the TOG pipeline using various VLMs and querying strategies. Our goal was to assess the effectiveness of integrating pre-trained grasp generation models with VLMs to select grasps that satisfy specific task requirements.
IV-A Experimental Setup
Our experiments were conducted in the ManiSkill simulator [17] using the Franka Emika Panda robotic arm equipped with a parallel-jaw gripper. We selected 29 objects from the YCB [5] and Google Scanned Objects (GSO) [6] datasets, ensuring diversity in shape, size, and functionality.
To evaluate task-oriented grasp generation, we defined three task categories: Grasping at a Specific Position, Tool Use, and Object Handover. Each task required the robot to grasp the object at a predefined position based on its intended function (see Table II).
Task Type Example Success Condition
Grasping at a Specific Position Grip the toy by its tail. The toy is grasped precisely at the specified position.
Tool Usage Hold the screwdriver by the handle for ease of use. The tool is held in a way that facilitates its intended function.
Object Handover Pass me the screwdriver. Hold the screwdriver by the shaft, leaving the handle free for easy transfer.
TABLE II: Three category of tasks using in simulation experiment
For evaluating task compliance, we initially attempted an automated assessment using Qwen-72B [28]. To verify its reliability, we manually reviewed a subset of results and compared them with Qwen-72B’s judgments. The model’s precision — defined as the agreement between its predictions and human evaluations — was 0.717. Given this suboptimal accuracy, we opted for manual evaluation following the predefined protocols in Table II.
We employed Contact GraspNet (CGN) [3] to generate candidate grasps for each object. CGN outputs grasps with associated quality scores, which were filtered using a motion planner to remove kinematically infeasible or unreachable grasps. We then followed the methodology described in Section III to generate a diverse set of grasp candidates. For the $k$-means algorithm, we set $k=3$, as it is robust across different VLMs.
Our experiments compared three Vision-Language Models (VLMs):
  • Qwen-VL [28]: An open-source VLM known for its strong language understanding and visual reasoning capabilities.
  • Molmo [29]: A VLM designed for robotic applications, particularly point-based interactions, making it well-suited for the PSI strategy.
  • Gemini-2.0-Pro [33]: A proprietary VLM demonstrating strong performance across diverse tasks.
We compared our proposed system against a baseline that selects the grasp with the highest confidence score from the grasp generation model.
@@ IV-A Experimental Setup
TABLE III: Experiment results of the proposed pipeline using different VLM query strategies for each VLM.
Figure 4: Comparison of grasp success, task success, and combined success across different evaluation metrics.
IV-B Evaluation Metrics
@@
IV-C Observations and Analysis
Table III summarizes our experimental results. Notably, the TOG system consistently outperforms the baseline system by a significant margin, demonstrating the effectiveness of the proposed pipeline for task-oriented grasp generation.
Is TOG effective for generating task-oriented grasps? Among the evaluated strategies, the Contact Point Generation (CPG) method achieves the best performance when combined with Gemini-2.0-Pro, as shown in Table III. However, CPG is less effective when used with weaker VLMs compared to other strategies. This result suggests that when leveraging a more powerful VLM, directly querying it to select the optimal grasp point based on the task description provides an accurate heuristic for selecting grasps generated by the CPG model. In contrast, when using a less powerful VLM, strategies that restrict the VLM to choose from a predefined set of clustered grasps perform better. Further analysis reveals that for weaker VLMs, the generated contact points are often unstable and sometimes misaligned with the object, leading to failures in grasp selection (see Figure 5 (c) and (d)). A key insight from our findings is that when using a more capable VLM, providing it with greater ”freedom” improves performance, whereas more restrictive approaches yield better results for less powerful models.
Does using a VLM affect grasp quality? We observe that the grasp success rate for methods incorporating the VLM is comparable to that of the baseline, indicating that integrating the VLM does not degrade grasp quality. Interestingly, the grasp success rate even increases when using a more powerful VLM in conjunction with the CPG method. We hypothesize that the VLM may implicitly favor more stable grasps, contributing to this improvement.
Which query strategy performs best? We observe that methods utilizing multiple images (GMI and CPMI) outperform those using a single image (GSI and CPSI), particularly when paired with Gemini-2.0-Pro. However, for less powerful VLMs, an increased number of images may introduce confusion, leading to a higher failure rate. Regarding query representation, we find that both strategies — visualizing the gripper shape (GSI and GMI) or the contact point (CPSI and CPMI) — perform comparably overall.
How to select the number of clusters $K$? We further investigate the effect of different $K$ values in our experiments. Figure 4 presents the results using Gemini-2.0-Pro. We compute the average score across all strategies that utilize clustering (GMI, CPMI, GSI, and GMI) and plot it against different $K$ values used in the $k$-means clustering algorithm. Notably, the task compliance rate increases as $K$ increases, whereas the grasp success rate decreases. This aligns with our observation that a higher $K$ allows the VLM greater freedom to select a grasp that best matches the task description. However, this increased freedom comes at a cost: greater diversity in clusters leads to a decline in grasp quality. High-quality grasps are typically supported by more neighboring grasps, which tend to form larger clusters. As a result, increasing $K$ leads to a trade-off between grasp diversity and grasp success rate.
In general, we find that selecting the optimal $K$ by maximizing the average silhouette score provides a good balance between grasp success rate and task compliance rate, as illustrated in Figure 4.
Refer to caption Refer to caption
(a) (b)
Refer to caption Refer to caption
(c) (d)
Figure 5: Common failure cases: (a) Insufficient grasp diversity due to clustering. (b) Ambiguous query image makes grasp identification difficult. (c) VLM selects an unstable grasp. (d) VLM assigns an incorrect contact point (the blue dot)
IV-D Failure Modes